OpenStudy (anonymous):

What are the solutions for the quadratic equation x2(to the second power) + 6 x = 16 ?

OpenStudy (anonymous):

easy

OpenStudy (tommo_lcfc):

Factorise. Found the solutions in 5 seconds! :) (Solutions are 8 and -2).

OpenStudy (anonymous):

(x+8)(x-2) x=-8 or 2

OpenStudy (anonymous):

you're wrong tommo

OpenStudy (anonymous):

gf your 5 secs

OpenStudy (anonymous):

how did you get 8 and 2

OpenStudy (anonymous):

Onc you have the factors of your quadratic, in this case x+8 and x-2, we set each one equal to zero and solve for x. What you are trying to do is find out when your quadratic function is equal to zero. This can happen in two ways, either x+8=0 or x-2=0. Solving each for x, we know our quadratic is equal to zero when x=2 or x=-8

OpenStudy (tommo_lcfc):

@Tai, I'm good at Maths, OK! I am doing a degree, so have to be! :) So I'm not wrong and you were kidding when you said I was, weren't you?

OpenStudy (anonymous):

what did you factor to get -8 and 2 ? if you factored 16 wouldnt it be 4 and 4 ???

OpenStudy (anonymous):

@tommo_lcfc x is 2 or -8...if you're doing a degree you've gotta reconsider re-taking highschool lol

OpenStudy (anonymous):

or maybe you're just tired?

OpenStudy (anonymous):

factorises to (x+8)(x-2) --> so you tell me what the values of x are lol.

OpenStudy (tommo_lcfc):

@ashleysmith 4 and 4 would not work as they cannot add up to 6!!! :)

OpenStudy (anonymous):

@tommo, yes, you are incorrect

OpenStudy (anonymous):

Hey don't judge him, everyone makes mistakes, there's nothing to be ashamed of, we're not robots ! ;)

OpenStudy (anonymous):

thanks guys !

OpenStudy (anonymous):

guys look...\[x^2+6x-16=0\] \[x^2+8x-2x-16=0\] \[x(x+8)-2(x+8)=0\] \[(x-2)(x+8)=0\] Hence: x=2 or x=-8

OpenStudy (anonymous):

and maybe i was a bit harsh, soz

OpenStudy (anonymous):

@ tommo_lcfc what kind of a degree are you taking?

OpenStudy (anonymous):

Another way to get the factors is to try to find two numbers that multiply to give -16 and add up to give +6. Some trial and error yields 8 and -2. Thus your factors are (x+8)(x-2)